OpenStudy (anonymous):

Can you get the domain and range of an inverse function without graphing it?

OpenStudy (anonymous):

you can, but it depends on your level of math.

OpenStudy (anonymous):

domain you can im sure.. probably not the range.

OpenStudy (anonymous):

what do you mean by level of math?

OpenStudy (anonymous):

for the domain, just see what values of x work for your problem. example.. if it is a linear equation.. y = 2x+1 then you know it can be all real values of x.

OpenStudy (anonymous):

I dont think they teach you how to find the range of an inverse function until calculus without making a graph and looking at it.

OpenStudy (anonymous):

is what i mean.

OpenStudy (anonymous):

oh. so you always need a graph to look for the range?

OpenStudy (anonymous):

yes, you have a graphic calculator? makes it much easier than drawing :)

OpenStudy (anonymous):

we can't use that. :|

OpenStudy (anonymous):

eww

OpenStudy (anonymous):

Can I ask a new question? Since you have answered my previous one.

OpenStudy (anonymous):

though once u find the domain and range for the inverse, the domain and range for f(x) is the opposite.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

There are some cases were the inverse of a function is not a function.

OpenStudy (anonymous):

well in our class my teacher gave a random domain in which he did not tell us how he got the domain, so can you explain how to give the right domain in those cases where the inverse is not a function?

OpenStudy (anonymous):

do you mean the inverse is not a one to one function?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

For example

OpenStudy (anonymous):

my teacher used this function: f(x)=(x-1)^2 -4

OpenStudy (anonymous):

then we got the inverse which is +- sqrt of x+4 +1 = y

OpenStudy (anonymous):

so what he did was to give a domain of x is greater than or equal to 1

OpenStudy (anonymous):

for the inverse

OpenStudy (anonymous):

so my question is, how did he get that domain? what was his basis for that domain?

OpenStudy (anonymous):

the domain has of \[\pm \sqrt{x+4+1}\] has to be greater than 0 because you cannot have a negative number under a radical sign

OpenStudy (anonymous):

when you are dealing with domain and functions.. you have to make sure that your answer does not break the rules... such as.. if you have a fraction.. x cannot = 0 in the denominator.. so you would leave that out of the domain.

OpenStudy (anonymous):

The original fxn was (x-1)^2 -4 and my teacher gave a domain of X is greater than or equal to 1.

OpenStudy (anonymous):

\[f(x)=(x-1)^2-4\] is a parabola that opens up, and has a vertex at \((1,-4)\)

OpenStudy (anonymous):

yay satellite is here! he is much smarter than I..

OpenStudy (anonymous):

if you make the domain \[x\geq 1\] then it is no longer a parabola, it is half of one, the part to the right of the vertex, and so it is increasing and one to one

OpenStudy (anonymous):

haha. :D so why is it confined into \[X \ge1\] only?

OpenStudy (anonymous):

so now it is a one to one function, and you can find the inverse

OpenStudy (anonymous):

why can't it be less than or equal to one?

OpenStudy (anonymous):

you get to pick the domain, so long as on that domain the function is one to one

OpenStudy (anonymous):

it could be lets solve that one

OpenStudy (anonymous):

so that means we can pick any domain we want? but how do we know if it is the right one?

OpenStudy (anonymous):

\[x=(y-1)^2-4\] \[x+4=(y-1)^2\] \[\pm\sqrt{x+4}=y-1\] \[y=\pm\sqrt{x+4}+1\] but since we know \(y\leq 1\) we get rid of the \(\pm\) and put \[y=-\sqrt{x+4}+1\]

OpenStudy (anonymous):

I'm kinda getting it now.

OpenStudy (anonymous):

there is no "right" domain you can pick any domain for which the original function is one to one in your case you could say domain is \(x\geq 1\) or you could say \(x\leq 1\) or for that matter you could say \(1<x<5\) because it is one to one there too

OpenStudy (anonymous):

sorry i hacked it up so much :)

OpenStudy (anonymous):

It's okay Nameless :D

OpenStudy (anonymous):

lets look at a picture of \[y=(x-1)^2-4\] http://www.wolframalpha.com/input/?i=y%3D%28x-1%29^2-4

OpenStudy (anonymous):

so when you say pick the domain... do you mean pick the numbers in the intervals between key numbers?

OpenStudy (anonymous):

you see it is not one to one, because it does not pass the horizontal line test but you can make it one to one by restricting the domain, in other words by taking a piece of the function

OpenStudy (anonymous):

So basically for this function: f(x)=(x-1)^2 -4 we can say that it's domain is: \[x \ge5 \]?

OpenStudy (anonymous):

you could say 'let \(f(x)=(x-1)^2-4\) for \(x\geq 1\) and then it will be one to one

OpenStudy (anonymous):

hey it is your function, you can do anything you like with it

OpenStudy (anonymous):

so basically that fxn can have lots of domain?

OpenStudy (anonymous):

you can restrict the domain in any way that you want

OpenStudy (anonymous):

that's cool :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks! Getting so confused there. :D

OpenStudy (anonymous):

hope that helped. the restriction of \(x\geq 1\) was not arbitrary, it was made as large as possible so that the function was one to one but \(x\leq1\) would work just as well, you would be taking the left side instead of the right side

OpenStudy (anonymous):

i am confused by the statement you can pick your domain

OpenStudy (anonymous):

Thank you so much satellite :D

OpenStudy (anonymous):

i can define a function any way i like

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i can say "let \(f(x)=x^2+2x\) for \(3<x<5\)

OpenStudy (anonymous):

so you can define to fit what you need?

OpenStudy (anonymous):

thus in this case only half of the parabola was needed to make this an inverse? so this is not an inverse function until you restrict the domain

OpenStudy (anonymous):

I guess that's what he's trying to say :D

OpenStudy (anonymous):

Thanks for your help as well Nameless :D

OpenStudy (anonymous):

np